OpenStudy (anonymous):

Find the first, fourth, and tenth terms of the arithmetic sequence described by the given rule. A(n)=1+(n-1)(-4.1) 1, –15.4, –40 –4.1, –15.4, –36.9 1, –11.3, –35.9 0, –12.3, –36.9

myininaya (myininaya):

To find first term replace n with 1 and evaluate

OpenStudy (anonymous):

would it be C @myininaya

myininaya (myininaya):

\[A(n)=1+(n-1)(-4.1) \\ A(1)=1+(1-1)(-4.1)=1+0=1 \\ A(2)=1+(2-1)(-4.1)=?\]

myininaya (myininaya):

I don't think that will output any of your choices

myininaya (myininaya):

are you sure you wrote your problem correctly?

myininaya (myininaya):

oh fourth term

myininaya (myininaya):

you can find the fourth term by replacing n with 4

myininaya (myininaya):

and then tenth term can be found by replacing n with 10

myininaya (myininaya):

so yeah that seems right